OpenStudy (anonymous):

The aspect ratio of a television screen is the length-to-width ratio. A standard definition television has an aspect ratio of 4:3. A high definition projection television had an aspect ratio of 16:9. Are the television screens similar rectangles? explain?

OpenStudy (anonymous):

Does 16/9 = 4/3?

OpenStudy (anonymous):

idk the answer is no: the ratio Length over length = 4/6=1/4 and the ratio width over width = 3/9=1/3;1/4\[\neq\]1/3

OpenStudy (anonymous):

but i dnt knw how they got that answer

OpenStudy (anonymous):

?

OpenStudy (anonymous):

The book is correct. Sure, well the ratio of each side must be equal for both ratios to be equal. You find the ratio of one length and compare it to the ratio of the other length. What this is doing is saying that for every single unit of the smaller one, the other one is going to be so much bigger. Like the length, for every inch that the smaller one has, the bigger one is 4 inches. The next part is to do the same thing and see if they are equal, but for every inch the smaller one has, the larger one only has 3 inches which means that the ratios aren't equal.

OpenStudy (anonymous):

oh !!! thank you very much... =D